A disintegrates into C by emitting 2 alpha and 3 beta particles, B converts into C by emitting 1 alpha and 5 beta particles.At t=0, A has 4N nuclei and B has N nuclei and their half lives are 1 and 2 minutes respectively,the time at which the rate of disintegration is equal.(at t=0,no. of nuclei of C=0).


 The answer is 6,how.

Grade:12

1 Answers

mounika gajawada
12 years ago

disintegration constants of

λA=ln2/1 and λB=ln2/2

as rates are equal; λANABNB

4N(e-{ln2/1}*t)*(ln2/1)=N(e-{ln2/2}*t)*(ln2/2)

→ 8(2-t)=(2-t/2)

t=6